OpenStudy (anonymous):

verify the identity: sin(π/2+x)=cosx

OpenStudy (zzr0ck3r):

sin ( pi/2 + x ) = sin(pi/2)cos(x) + cos(pi/2)sin(x) = 1*cos(x) + 0

OpenStudy (anonymous):

how did you get to that point?

OpenStudy (zzr0ck3r):

sin(a+b) = sin(a)cos(b) + cos(a)sin(b)

OpenStudy (anonymous):

Addition identities. They're pretty great.

OpenStudy (zzr0ck3r):

sin(A+B)=sin A cos B + cos A sin B sin(A-B)=sin A cos B - cos A sin B cos(A+B)=cos A cos B - sin A sin B cos(A-B)=cos A cos B + sin A sin B

OpenStudy (anonymous):

oh yes, i remember those formulas! thanks for the refresher ;)

OpenStudy (zzr0ck3r):

np, I don't remember them I google them:)

OpenStudy (zzr0ck3r):

well ok sin(a+b) I remembered..

OpenStudy (anonymous):

hahaha @zzr0ck3r ok, thanks anyways

OpenStudy (anonymous):

\[\sin{\pi\over2}\cos x+\cos{\pi\over2}\sin x\]\[(1)(\cos x)+0(\sin x)\]\[\cos x+0\]\[\sin({\pi\over2}+x)=\cos x\]

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

that's pretty:)

OpenStudy (anonymous):

what is? lol

OpenStudy (anonymous):

i have no idea..maybe your beautiful math response...its neatly typed haha

OpenStudy (anonymous):

lol, hope i've been helpful! :)

OpenStudy (anonymous):

yes, thanks! @yummydum

OpenStudy (anonymous):

Anytime! :)